OpenStudy (anonymous):

multiply and simplify v^2-5v-36/v^2-8v+16 x v-4/v-9

OpenStudy (anonymous):

I've got a few screws missing too...hence why i cannot do math LOL

OpenStudy (hoblos):

z+4 = --- z-4

OpenStudy (hoblos):

do u need explanation

OpenStudy (anonymous):

yes please :D

OpenStudy (hoblos):

v^2 -5v -36 = (z-9)(z+4) v^2-8v+16 = (z-4)^2 now replace them

OpenStudy (hoblos):

get it

OpenStudy (anonymous):

no im still confused as pellet

OpenStudy (hoblos):

mmm it seems that u have problems with factoring don't you?